OpenStudy (anonymous):

write in standard form an equation of the line passing through the given point and slope. slope=8/7;(5,-3) PLEASE SHOW WORK

OpenStudy (anonymous):

\[(y-y _{1}) = m(x-x _{1})\] now replace\[y _{1} =-3\] and \[x _{1} =5\] and \[m =8/7\]

OpenStudy (precal):

 use this

OpenStudy (precal):